OpenStudy (fatimah66):

What is one of the solutions to the following system of equations? y2 + x2 = 65 y + x = 7

OpenStudy (opti):

Use substitution

OpenStudy (radar):

Express y in terms of x using the 2nd equation. y = 7 - x Now determine y^2 using (7 - x)^2 getting x^2 - 14y + 49 Substitute this for Y^2 in the first equation getting: x^2 -14y + 49 + x^2 = 65 Now simplify and solve the quadratic by factoring.

OpenStudy (radar):

As opti suggested "use substitution"

OpenStudy (radar):

@fatimah66 Can you continue the problem getting the solution?

OpenStudy (radar):

If not, I can continue to the point where you can take the hand-off.

OpenStudy (radar):

Just say the word.

OpenStudy (radar):

next step. simplify by combining similar terms getting: 2x^2 - 14x + 49 = 65 2x^2 -14x - 16 = 0 Divide thru by 2 getting x^2 - 7x - 8 = 0 Now solve that quadratic (suggest factoring). Good luck with your studies.